OpenStudy (anonymous):

A segment is divided into two parts having lengths in the ratio of 5:3. If the difference between the length of the parts is 6", find the length of the longest part. 3 9 15 24

OpenStudy (anonymous):

lets the length of both parts are x and y... than x/y=5/3....(1) and x-y=6... x=6+y... putting the value of x in equation 1... 6+y/y=5/3... 18+3y=5y... 18=5y-3y.. 2y=18... y=9 cm... x=6+y... x=6+9.. x=15... longest part length is 15 cm.